# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ 2 Problem Description
2.1 Differentially Private Histogram
@@
We require that both the output $\tilde{h}$ and the running time of the algorithm reveal only a limited amount of information about the input. The privacy of the output is formally captured by the notion of differential privacy.
Neighboring Datasets. Two datasets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ are said to be neighboring, denoted $\mathcal{X}\sim\mathcal{X}^{\prime}$, if they differ in at most one participant’s data. There are two standard notions of neighboring datasets:
  • $\triangleright$
    Replacement Neighboring: $\mathcal{X}$ and $\mathcal{X}^{\prime}$ have the same size and differ in exactly one participant’s data; that is, there exists a unique $u\in[n]$ such that $x^{({u})}\neq x^{\prime\,({u})}$ and $x^{({i})}=x^{\prime\,({i})}$ for all $i\neq u$.
  • $\triangleright$
    Add/Remove Neighboring: $\mathcal{X}$ and $\mathcal{X}^{\prime}$ differ by the addition or removal of a single entry; that is, $\mathcal{X}^{\prime}=\mathcal{X}\cup\{x^{({n+1})}\}$ or $\mathcal{X}^{\prime}=\mathcal{X}\setminus\{x^{({i})}\}$ for some $i\in[n]$.
Accordingly, two histograms $\vec{h},\vec{h}^{\prime}\in\mathbb{N}^{d}$ are called neighboring if they are induced by neighboring datasets under one of the above definitions. Throughout this paper, we primarily adopt the replacement neighboring model. An extension of our algorithm to the add/remove model is discussed in Section 4.3.
@@
Let $\varepsilon\in\mathbb{R}_{+}$, $\delta\in[0,1]$, and let $(\mathcal{Z},\mathcal{F})$ be a measurable space. Two probability measures $\mu$ and $\nu$ on this space are said to be ${({\varepsilon,\delta})}$-indistinguishable if
$e^{-\varepsilon}\cdot{({\mu(E)-\delta})}\leq\nu(E)\leq e^{\varepsilon}\cdot\mu(E)+\delta,\quad\forall E\in\mathcal{F}.$ (1)
@@
Given $\varepsilon\in\mathbb{R}_{+}$, $\delta\in[0,1]$, a randomized algorithm $\mathcal{M}:\mathcal{Y}\rightarrow\mathbb{N}^{d}$ is called ${({\varepsilon,\delta})}$-differentially private (DP), if for every $\mathcal{X},\mathcal{X}^{\prime}\in\mathcal{Y}$ such that $\mathcal{X}\sim\mathcal{X}^{\prime}$, $\mathcal{M}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X}^{\prime})$ are ${({\varepsilon,\delta})}$-indistinguishable.
An algorithm $\mathcal{M}$ is said to be pure differentially private, or simply $\varepsilon$-DP, if it satisfies $(\varepsilon,0)$-differential privacy. If instead $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy for some $\delta>0$, it is referred to as approximate differentially private. For a histogram release algorithm, the input domain $\mathcal{Y}$ depends on the chosen neighboring model. Under the replacement model, $\mathcal{Y}=\mathcal{D}^{n}$, whereas under the add/remove model, $\mathcal{Y}=\bigcup_{k\in\mathbb{N}_{+}}\mathcal{D}^{k}$.
In the add/remove model, there is no fixed upper bound on the dataset size, so it is also referred to as the unbounded setting. A commonly studied relaxation is the upper-bounded setting, where a public upper bound on the dataset size is assumed.
@@
Time Obliviousness. The privacy of the running time is captured with by an extended notion differential privacy. Formally, we require the joint distribution of the algorithm’s output and running time to satisfy differential privacy.
Definition 2.3 (${({\varepsilon,\delta})}$-DP Time-Oblivious Algorithm (Dov et al., 2023)).
A straightforward approach to designing DP time-oblivious algorithms is to establish a deterministic runtime bound—preferably polynomial in the input size—and pad execution up to this bound, thereby preventing timing side-channel leaks.
@@
We evaluate utility using the following standard accuracy metrics.
@@
Apart from obtaining $(\alpha,\beta)$-simultaneous accuracy bounds for $\tilde{h}$, we also consider bounds on the expected error $\operatorname*{\mathbb{E}}\left[\lVert{\tilde{h}-\vec{h}}\rVert_{\infty}\right]$. In fact, the latter can be derived from the former using standard integration techniques; see Section 4 for details. Some prior work also considers a per-query accuracy guarantee.
Definition 2.5 ($(\alpha,\beta)$-Accurate Estimator).
A histogram $\tilde{h}$ achieves $(\alpha,\beta)$ per-query accuracy if, for each $i\in[d]$, $\tilde{h}{\left[{i}\right]}$ is an $(\alpha,\beta)$-accurate estimator of $\vec{h}{\left[{i}\right]}$. In the sparse histogram setting, these two notions coincide. Balcer and Vadhan (2019) show that, for any $\beta\in(0,1/2]$, an $\varepsilon$-DP histogram $\tilde{h}$ with $\lVert{\tilde{h}}\rVert_{0}\leq n^{\prime}$ cannot achieve $(\alpha,\beta)$ per-query accuracy for $\alpha\in\Omega{\big{(}{\min\{{\frac{1}{\varepsilon}\ln\frac{d}{n^{\prime}\cdot\beta},\,n}\}}\big{)}},$ which matches the simultaneous accuracy upper bounds obtained by our algorithm in Section 4.
There exist other private succinct representations of $\vec{h}$ beyond the sparse histogram, which can bypass the per-query accuracy lower bound (Balcer and Vadhan, 2019; Aumüller et al., 2021; Lolck and Pagh, 2024). However, these representations typically incur higher computational cost when estimating $\vec{h}{\left[{i}\right]}$ for a given $i\in[d]$. Moreover, they are inadequate for solving the heavy hitters task—i.e., identifying all $i\in[d]$ such that $\vec{h}{\left[{i}\right]}\in\Omega\left(\frac{1}{\varepsilon}\cdot\ln d\right)$—unless one performs $\tilde{O}(d)$ queries to the frequency of every domain element. Thus, this line of work is orthogonal to our focus on sparse histograms. We elaborate on this distinction in the related work section (Section 6).
3 Preliminaries
In this section, we define several probability distributions used throughout the paper, and review two building blocks for constructing our algorithm: an array-based algorithm for sampling from distribution with finite support, and a framework of converting approximate DP into pure DP algorithms.
3.1 Probability Distributions
@@
3.2 Alias Method
The Alias method (Walker, 1977), denoted by ${\mathcal{M}_{\mathbb{A}lias}}$, is an efficient array-based sampling algorithm that, given a distribution $\mu=\{{p_{0},\ldots,p_{m-1}}\}$ over $\mathbb{N}_{<{m}}$, has $O(m)$ initialization time and memory usage, and then draw samples from $\mu$ in $O(1)$ time. The pseudocode is given in Algorithm 1.
During the initialization phase, the algorithm constructs two arrays $\vec{a}$ and $\vec{b}$ of length $\bar{m}\geq m$. The array $\vec{a}\in\mathbb{N}_{<{m}}^{\bar{m}}$ is called the alias array, and $\vec{b}\in[0,1]^{\bar{m}}$ stores Bernoulli probabilities.
Sampling Procedure (${\mathcal{M}_{\mathbb{A}lias}.\textsc{\small Sample}}$): To sample from $\mu$, the algorithm proceeds as follows: (1) sample a uniform random index $I\in\mathbb{N}_{<{\bar{m}}}$, (2) sample a Bernoulli random variable $B$ that equals $1$ probability $\vec{b}[I]$, (3) if $B=1$, return $I$; otherwise, return $\vec{a}[I]$. That is why $\vec{a}$ is called the alias array: it specifies the alternative index to return when the Bernoulli test fails.
Proposition 3.1 ((Walker, 1977)).
For completeness, we include the pseudocode as well as the proof of the construction of $\vec{a}$ and $\vec{b}$ in Appendix B. It is worth noting that Walker (1977) assumes real arithmetic, which is infeasible in the word-RAM model used in this paper. In Section 5.2, we relax this assumption and apply the Alias method to approximately sample from a discrete distribution.
Algorithm 1 Alias Method ${\mathcal{M}_{\mathbb{A}lias}}$
1: Procedure: Initialization
2: Input: $m\in\mathbb{N}^{+}$, $p_{0},\ldots,p_{m-1}\in{({0,1})}$
3:Construct arrays $\vec{a}$ and $\vec{b}$ as in Proposition 3.1 (Walker, 1977);
4:See pseudocode in Appendix B for completeness.
5: Procedure: Sample
6: Output: Random variable $Z\in\mathbb{N}_{<{m}}$
7:$I\leftarrow\operatorname{\mathbb{U}niform}\left({\mathbb{N}_{<{\bar{m}}}}\right)$, $B\leftarrow\operatorname{\mathbb{B}ernoulli}\left({\vec{b}[I]}\right)$
8:if $B=1$ then return $Z\leftarrow I$
9:if $B=0$ then return $Z\leftarrow\vec{a}[I]$
3.3 Purification
Balcer and Vadhan (2019) proved that, given an algorithm $\mathcal{M}^{\prime}:\mathcal{Y}\rightarrow\mathcal{Z}$ with finite output domain, if there exists a pure DP algorithm $\mathcal{M}:\mathcal{Y}\rightarrow\mathcal{Z}$ such that the total variation distance between the output distributions of $\mathcal{M}^{\prime}$ and $\mathcal{M}$ is sufficiently small, then one can convert $\mathcal{M}^{\prime}$ into a pure DP algorithm by mixing its output distribution with a uniform distribution over $\mathcal{Z}$. The pseudocode is provided in Algorithm 2, and the formal statement is in Proposition 3.2.
@@ 3.3 Purification
3:With probability $1-\gamma$, output $\mathcal{M}^{\prime}(y)$
4:Otherwise, output $Z\sim\operatorname{\mathbb{U}niform}\left({\mathcal{Z}}\right)$
Proposition 3.2 (Balcer and Vadhan (2019)).
@@
    Running time $O{\left({\frac{1}{\omega}\cdot\log\frac{1}{\gamma}+\frac{1}{\omega}\cdot\log\frac{1}{\lvert{\mathcal{Z}}\rvert}}\right)}+T_{{\mathcal{M}^{\prime}}}$, where $T_{{\mathcal{M}^{\prime}}}$ is the running time of $\mathcal{M}^{\prime}$.
4 Differentially Private Sparse Histograms
@@
Given integers $n,d,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma}\in\mathbb{N}_{+}$ that fit in a constant number of machine words, where $\gamma\doteq a_{\gamma}/b_{\gamma}\in{({1/n^{O(1)},1})}$, and a histogram $\vec{h}\in\mathbb{N}_{\leq{n}}^{d}$ such that $\lVert{\vec{h}}\rVert_{1}\leq n$, there exists an algorithm, denoted by $\mathcal{M}_{\textsc{\scriptsize Hist}}$, that outputs a sparse approximation $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$ of $\vec{h}$, satisfying $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and
@@
    Privacy Guarantee: $\mathcal{M}_{\textsc{\scriptsize Hist}}$ satisfies $2\varepsilon$-differential privacy, where $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}\in\mathbb{Q}_{+}$.
@@
    Running Time: $\mathcal{M}_{\textsc{\scriptsize Hist}}$ runs deterministically in $O(n\ln\ln d)$ time.
Roadmap. The remainder of this section is organized as follows. We begin by discussing Theorem 4.1 and its connection to the informal version stated in the introduction (Theorem 1.1). Section 4.1 presents the key ideas behind our algorithm, along with its pseudocode. Section 4.2 provides the proof of Theorem 4.1. Since this algorithm assumes the replacement neighboring model, Section 4.3 shows how to extend it to the add/remove model.
A comprehensive comparison of our algorithm with prior work is deferred to Section 6. A prototype extension to the MPC setting is presented in Appendix D.
[…]
Also, in what follows, we assume $d\geq 10n$ to focus on the sparse regime where our algorithm is most relevant. Otherwise, when $d<10n$, one can simply release the entire histogram $\vec{h}$ after perturbing each count using the algorithm in Theorem 1.2, which yields an algorithm with $O(n)$ running time.
4.1 Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
To motivate our algorithm, we begin by reviewing the Filter algorithm for releasing pure-DP sparse histograms (Cormode et al., 2012), which lacks a known efficient time-oblivious implementation. Given an input histogram $\vec{h}\in\mathbb{N}_{\leq{n}}^{d}$ with $\lVert{\vec{h}}\rVert_{1}=n$, the Filter algorithm produces an output $\tilde{h}$ distributed identically to the output of the following procedure: (1) Add independent $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ noise to each entry of $\vec{h}$. (2) Retain only the entries whose noisy values exceed a threshold $\tau\in\Theta\left(\frac{1}{\varepsilon}\cdot\ln d\right)$, chosen so that the expected number of outputs is $O(n)$.
@@ 4.1 Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
    Sample the number $X$ of zero entries exceeding $\tau$ from a binomial distribution $\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})$, where $\operatorname{\textsc{supp}}(\vec{h})$ denotes the support of $\vec{h}$.
@@
    Sample a random subset of size $X$ from $[d]\setminus\operatorname{\textsc{supp}}(\vec{h})$.
@@
Though Filter has a simple description, implementing it properly is not straightforward. The first issue is that sampling from $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ exactly is vulnerable to timing attacks. To address this, one can replace $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ with a noise distribution with similar properties that has deterministic sampling time, such as the one presented in Theorem 1.2. In addition, one can avoid sampling from the conditional distribution in the last step by instead using fresh noises, at the slight cost of a factor-$2$ degradation in utility.
The real challenge lies in sampling from the distribution $\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})$. Even writing down the exact probabilities of this binomial distribution requires exponential space. A natural idea is to approximately sample: generate a random variable $X$ such that $\mathds{TV}({X},\,{\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})})\leq\delta$ for some small $\delta$. This leads to an algorithm that satisfies approximate DP. To recover pure DP, one could apply the purification technique (Proposition 3.2), which mixes the final output distribution of Filter with the uniform distribution over the output space to smooth out the $\delta$ gap. However, Proposition 3.2 requires $\delta\in O(1/2^{d})$, since the output of Filter may be any subset of $[d]$. For such a small $\delta$, we do not know how to efficiently sample $X$ satisfying $\mathds{TV}({X},\,{\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})})\leq\delta$.
@@
2: parameter $\tau\in\mathbb{N}_{+}$triangleright as defined in Equation 2
3: histogram $\vec{h}\in\mathbb{N}_{\leq{n}}^{d}$ with $\lVert{\vec{h}}\rVert_{1}=n$
4: Output: histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$
@@
6:     $\hat{h}{\left[{i}\right]}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}{({\vec{h}{\left[{i}\right]}})}$triangleright $\hat{h}{\left[{i}\right]}\approx\text{clamp}{[{{\vec{h}{\left[{i}\right]}+\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)},{0},{n}}]}$
7:$I_{1}\leftarrow\{i\in\operatorname{\textsc{supp}}{({\vec{h}})}:\hat{h}{\left[{i}\right]}\geq\tau\}$
8:$I_{2}\leftarrow$ a uniform random subset in $\binom{[d]\setminus I_{1}}{n+k-\lvert{I_{1}}\rvert}$
9:$I\leftarrow I_{1}\cup I_{2}$
@@
11:     $\tilde{h}{\left[{i}\right]}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}{({\vec{h}{\left[{i}\right]}})}$triangleright $\tilde{h}{\left[{i}\right]}\approx\text{clamp}{[{{\vec{h}{\left[{i}\right]}+\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)},{0},{n}}]}$
12:return $\tilde{h}$triangleright where $\tilde{h}[i]=0$ for all $i\notin I$
@@
Our goal is to design a simpler privacy blanket that satisfies both requirements. The pseudocode for our algorithm is provided in Algorithm 3.

### v2

@@ 2. Problem Description
2.1. Differentially Private Histogram
@@
We require that the output $\tilde{h}$ reveal only a limited amount of information about the input. The privacy of the output is formally captured by the notion of differential privacy.
Neighboring Datasets.
Consistent with Balcer and Vadhan (2019), we adopt the replacement neighboring model [^3] Another commonly used notion is the add/remove neighboring model. However, under this model, no optimal-error private frequency histogram algorithm can run in deterministic time. A formal definition of the model and further discussion appear in Appendix E. This limitation is precisely why Balcer and Vadhan (2019) restrict their study to the replacement model, where the question of breaking the quadratic running-time barrier was originally posed. , where two datasets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of equal size are neighboring, denoted $\mathcal{X}\sim\mathcal{X}^{\prime}$, if they differ in exactly one participant’s data; that is, there exists a unique $u\in[n]$ such that $x^{({u})}\neq x^{\prime\,({u})}$ and $x^{({i})}=x^{\prime\,({i})}$ for all $i\neq u$. Accordingly, two histograms $\vec{h},\vec{h}^{\prime}\in\mathbb{N}^{d}$ are called neighboring if they are induced by neighboring datasets under one of the above definitions.
Private Algorithm.
@@
Let $\varepsilon\in\mathbb{R}_{+}$, $\delta\in[0,1]$, and let $(\mathcal{Z},\mathcal{F})$ be a measurable space. Two probability measures $\mu$ and $\nu$ on this space are $({\varepsilon,\delta})$}-indistinguishable if
(1) $e^{-\varepsilon}\cdot{({\mu(E)-\delta})}\leq\nu(E)\leq e^{\varepsilon}\cdot\mu(E)+\delta,\quad\forall E\in\mathcal{F}.$
@@
Given $\varepsilon\in\mathbb{R}_{+}$, $\delta\in[0,1]$, a randomized algorithm $\mathcal{M}:\mathcal{D}^{n}\rightarrow\mathbb{N}^{d}$ is called $({\varepsilon,\delta})$}-differentially private (DP), if for every $\mathcal{X},\mathcal{X}^{\prime}\in\mathcal{Y}$ such that $\mathcal{X}\sim\mathcal{X}^{\prime}$, $\mathcal{M}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X}^{\prime})$ are $({\varepsilon,\delta})$}-indistinguishable.
An algorithm $\mathcal{M}$ is said to be pure differentially private, or simply $\varepsilon$-DP, if it satisfies $(\varepsilon,0)$-differential privacy. If instead $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy for some $\delta>0$, it is referred to as approximate differentially private.
@@
We evaluate utility using the $\ell_{\infty}$ error (Balcer and Vadhan, 2019).
@@
In addition to $(\alpha,\beta)$-simultaneous guarantees for $\tilde{h}$, we also provide bounds on the expected error $\mathbb{E}\big[\lVert{\tilde{h}-\vec{h}}\rVert_{\infty}\big]$; the latter follows from the former using standard integration techniques (see Section˜4 for details).
3. Preliminaries
In this section, we define several probability distributions used throughout the paper, and review a framework of converting approximate DP into pure DP algorithms.
3.1. Probability Distributions
@@
3.2. Purification
Balcer and Vadhan (2019) showed that if an algorithm $\mathcal{M}^{\prime}:\mathcal{Y}\to\mathcal{Z}$ with finite output domain is close in total variation distance to some pure DP algorithm $\mathcal{M}:\mathcal{Y}\to\mathcal{Z}$, then one can convert $\mathcal{M}^{\prime}$ into a pure DP algorithm by mixing its output with the uniform distribution over $\mathcal{Z}$. The purification procedure is given in Algorithm˜1: with probability $1-\gamma$, the algorithm outputs $\mathcal{M}^{\prime}(y)$, and otherwise outputs a uniformly random $Z\in\mathcal{Z}$. The formal guarantee is stated in Proposition˜3.3.
@@ 3.2. Purification
[…]
4:Output ${({1-B})}\cdot\mathcal{M}^{\prime}(y)+B\cdot Z$, where $Z\sim\operatorname{\mathbb{U}niform}\left({\mathcal{Z}}\right)$.
Proposition 3.0 ((Balcer and Vadhan, 2019)).
@@
    Running time $O{\left({\frac{1}{\omega}\cdot\log\frac{1}{\gamma}+\frac{1}{\omega}\cdot\log\lvert{\mathcal{Z}}\rvert}\right)}+T_{{\mathcal{M}^{\prime}}}$, where $T_{{\mathcal{M}^{\prime}}}$ is the running time of $\mathcal{M}^{\prime}$.
4. Differentially Private Sparse Histograms
@@
Given integers $n,d,k,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma}\in\mathbb{N}_{+}$ that fit in a constant number of machine words, where $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}\in\mathbb{Q}_{+}$ and $\gamma\doteq a_{\gamma}/b_{\gamma}\in{({1/n^{O(1)},1})}$, and a dataset $\mathcal{X}\doteq\{{x^{({1})},\ldots,x^{({n})}}\}\in[d]^{n}$, there exists an algorithm, denoted by $\mathcal{M}_{\textsc{\scriptsize Hist}}$, that outputs a histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$ approximating the original histogram $\vec{h}$ of $\mathcal{X}$, satisfying $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and
@@
    Privacy Guarantee: $\mathcal{M}_{\textsc{\scriptsize Hist}}$ satisfies $\max{\Big({\frac{3\varepsilon}{2}+\frac{1}{k+1},2\varepsilon}\Big)}$-differential privacy.
@@
    Running Time: $\mathcal{M}_{\textsc{\scriptsize Hist}}$ runs deterministically in $O(n+k)$ time.
Remark.
The utility guarantee of Theorem˜4.1 is an $\ell_{\infty}$ bound and is asymptotically optimal, matching the lower bound $\Omega{({\min\{{\tfrac{1}{\varepsilon}\ln d,n}\}})}$ of Balcer and Vadhan (2019). The bound provided here is a high-probability one, in contrast to the expected one stated in Theorem˜1.1. The latter can be recovered from the former by the standard identity $\mathbb{E}[X]=\int_{0}^{\infty}\Pr[X\geq t]\,dt$ for a nonnegative random variable $X$.
Setting $k=2/\varepsilon$ yields a privacy guarantee of $2\varepsilon$-DP and running time $O(n+1/\varepsilon)$, which simplifies to $O(n)$ in the meaningful regime $n\geq 1/\varepsilon$. When $n<1/\varepsilon$, simply returning the all-zero frequency vector incurs error at most $n<1/\varepsilon$, matching the lower bound.
Also, in what follows, we assume $d\geq 10n$ to focus on the sparse regime where our algorithm is most relevant. Otherwise, when $d<10n$, one can simply release the entire histogram $\vec{h}$ after perturbing each count using the algorithm in Theorem˜1.4, which yields an algorithm with $O(n)$ running time.
Roadmap.
The remainder of this section is organized as follows. We begin by discussing Theorem˜4.1 and its connection to the informal version stated in the introduction (Theorem˜1.1). Section˜4.1 presents the key ideas behind our algorithm, along with its pseudocode. Section˜4.2 provides the proof of Theorem˜4.1.
We present an MPC-friendly prototype circuit implementation of our algorithm in Appendix˜D. A detailed comparison with prior work is deferred to Appendix˜F.
4.1. Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
To motivate our algorithm, we begin by reviewing the Filter algorithm for releasing pure-DP sparse histograms (Cormode et al., 2012), which lacks a known efficient time-oblivious implementation. Given a histogram $\vec{h}\in\mathbb{N}_{\leq{n}}^{d}$ with $\lVert{\vec{h}}\rVert_{1}=n$, the Filter algorithm produces an output $\tilde{h}$ distributed identically to the output of the following procedure: (1) Add independent $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ noise to each entry of $\vec{h}$. (2) Retain only the entries whose noisy values exceed a threshold $\tau\in\Theta\left(\frac{1}{\varepsilon}\cdot\ln d\right)$, chosen so that the expected number of outputs is $O(n)$.
@@ 4.1. Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
    Sample the number $X$ of zero entries exceeding $\tau$ from a binomial distribution $\operatorname{\mathbb{B}in}(d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert,q_{\tau})$ where $\operatorname{\textsc{supp}}(\vec{h})$ denotes the support of $\vec{h}$.
@@
    Sample a random subset of size $X$ from $d]\setminus\operatorname{\textsc{supp}}(\vec{h})$.
@@
Though Filter has a simple description, implementing it properly is not straightforward. The first issue is that sampling from $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ exactly is vulnerable to timing attacks. To address this, one can replace $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ with a noise distribution with similar properties that has deterministic sampling time, such as the one presented in Theorem˜1.4. In addition, one can avoid sampling from the conditional distribution in the last step by instead using fresh noises, at the slight cost of a factor-$2$ degradation in utility.
The real challenge lies in sampling from the distribution $\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})$. Even writing down the exact probabilities of this binomial distribution requires exponential space. A natural idea is to approximately sample: generate a random variable $X$ such that $\mathds{TV}({X},\,{\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})})$ $\leq\delta$ for some small $\delta$. This leads to an algorithm that satisfies approximate DP. To recover pure DP, one could apply the purification technique (Proposition˜3.3), which mixes the final output distribution of Filter with the uniform distribution over the output space to smooth out the $\delta$ gap. However, Proposition˜3.3 requires $\delta\in O(1/2^{d})$, since the output of Filter may be any subset of $[d]$. For such a small $\delta$, we do not know how to efficiently sample $X$ satisfying $\mathds{TV}({X},\,{\operatorname{\mathbb{B}in}({d-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert},{q_{\tau}})})\leq\delta$.
@@
2: dataset $\mathcal{X}\doteq\{{x^{({1})},\ldots,x^{({n})}}\}\in[d]^{n}$
3: Output: histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$
4:construct histogram $\vec{h}\in\mathbb{N}_{\leq{n}}^{d}$ based on $\mathcal{X}$
@@
6:  $\hat{h}{\left[{i}\right]}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}{({\vec{h}{\left[{i}\right]}})}$ $\triangleright$ $\hat{h}{\left[{i}\right]}\approx\text{clamp}{[{{\vec{h}{\left[{i}\right]}+\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)},{0},{n}}]}$
[…]
@@
11:  $\tilde{h}{\left[{i}\right]}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}{({\vec{h}{\left[{i}\right]}})}$ $\triangleright$ $\tilde{h}{\left[{i}\right]}\approx\text{clamp}{[{{\vec{h}{\left[{i}\right]}+\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)},{0},{n}}]}$
12:return $\tilde{h} \triangleright$ where $\tilde{h}[i]=0$ for all $i\notin I$
@@
Our goal is to design a simpler privacy blanket that satisfies both requirements. The pseudocode for our algorithm is provided in Algorithm˜2. The noise sampler $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ used in Algorithm˜2 is described in Theorem˜1.4. This mechanism produces noise with the same privacy guarantees and nearly identical tail bounds as discrete Laplace noise, but clamped to the range $[0,n]$. For ease of exposition, it is helpful to conceptually treat these samples as clamped discrete Laplace noise.